OpenStudy (anonymous):

If g(x) = 6 + x + e^x, find g^-1of (7).

OpenStudy (anonymous):

try x=0 see if \[g(0)=7\] then you will have your answer

OpenStudy (anonymous):

you really have to guess this one. no computing it using algebra

OpenStudy (anonymous):

ln(x-6)/2 gives me 0/2 or zero is this right

OpenStudy (anonymous):

Plug g(x)=7, and solve for x.

OpenStudy (anonymous):

that is not the inverse. you would have to solve \[x=6+y+e^y\] for y. good luck!

OpenStudy (anonymous):

yes but for this one you have to guess. you check that \[g(0)=6+0+e^0=6+1=7\] so \[g^{-1}(7)=0\]

OpenStudy (anonymous):

don't try to find \[g^{-1}(x)\] using algebra

OpenStudy (anonymous):

hello anwar. seen quite a spate of wrong answers here tonight

OpenStudy (anonymous):

not from you of course

OpenStudy (anonymous):

Yeah, Plugging \(7\) for g(x) gives: \(x+e^x=1\), then \(x=0\) is the only solution. Hence \(g^{-1}(7)=0\).

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

exactly but you solved the last one by inspection yes?

OpenStudy (anonymous):

Yes! But, it's not hard to see it. It can be easily solved by graphing as well.

OpenStudy (anonymous):

And hello satellite :)